A firm has a total cost function equal to:


TC(q)=20+4q+0.1q^2



And faces a demand curve given by


P(q)= 15-0.3Q


What is the profit maximising production quantity, q, for this firm?
1
Expert's answer
2020-11-20T09:27:23-0500

To determine the profit maximizing production quantity the steps below should be followed;

i) Determine the Marginal Cost from the the Total Cost function.

"TC_{(q)}=20+4q+0.1q^2"

Differentiate the Total cost in respect to quantity to obtain the marginal cost.

"Therefore\\
 \frac{\Delta\ TC}{\Delta\ q}=4+0.2q\\
MC=4+0.2q"

ii) Determine the Total revenue from the price function.

"Total\ Revue(TR)=Price(p)\times\ quantity(q)\\
TR=q(15-0.3q)\\
=15q-0.3q^2"

Differentiate the function to obtain the Marginal Revenue.

"MR=\frac{\Delta\ TR}{\Delta\ q}=15-0.6q"

iii) To determine the maximum production quantity then, the Marginal Revenue"(MR)" =Marginal Cost"(MC)"

"Therefore\ 15-0.6q=4+0.2q\\
q=13.5"